OpenStudy (unklerhaukus):

Prove Euler's Equation \[x^2\frac{\text d^2y}{\text d{x^2}}+\alpha x\frac{\text dy}{\text dx}+by=f(x)\]\[\Downarrow\qquad\Downarrow\]\[\frac{\text d^2y}{\text d{t^2}}+(\alpha-1)\frac {\text dy}{\text dt}+by=g(t)\]

OpenStudy (unklerhaukus):

let \(\ln x=t\) \[x=e^t\]\[\frac{\text dx}{\text dt}=e^t\qquad \frac{\text d^2x}{\text dt^2}=e^t\]

OpenStudy (anonymous):

\[\frac{dy}{dx}=\frac{dy}{dt} \frac{dt}{dx}=e^{-t} \frac{dy}{dt}\] \[\frac{d^2y}{dx^2}=\frac{d}{dx}(\frac{dy}{dx})=\frac{d}{dt}(\frac{dy}{dx})\frac{dt}{dx}=e^{-t} \frac{d}{dt}(e^{-t} \frac{dy}{dt})=e^{-2t}(\frac{d^2y}{dt^2})-e^{-t} \frac{dy}{dt}\]

OpenStudy (unklerhaukus):

//

OpenStudy (anonymous):

plug in the original equation \[e^{2t} (e^{-2t} \frac{d^2y}{dt^2}-e^{-2t}\frac{dy}{dt})+\alpha e^{t} (e^{-t} \frac{dy}{dt}) +by=g(t)\ \\ \frac{d^2y}{dt^2}+(\alpha-1)\frac{dy}{dt}+by=g(t)\]

OpenStudy (anonymous):

there is a little typo in my first reply second line last term e^(-2t)*****

OpenStudy (unklerhaukus):

Fantastic work mukushala ,

OpenStudy (anonymous):

tnx my friend

OpenStudy (unklerhaukus):

but the typo your referring to is in the last term right? \[\frac{d^2y}{dx^2}=\cdots=e^{-2t}(\frac{d^2y}{dt^2})-e^{-2t} \frac{dy}{dt}\]

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

\[\frac{d^2y}{dx^2}=\frac{d}{dx}\left(\frac{dy}{dx}\right)={\color\blue{\frac{d}{dt}\left(\frac{dy}{dx}\right)\frac{dt}{dx}}}\color\red=e^{-t} \frac{d}{dt}\left(e^{-t} \frac{dy}{dt}\right)=e^{-2t}\left(\frac{d^2y}{dt^2}\right)-e^{-2t} \frac{dy}{dt}\]

OpenStudy (unklerhaukus):

i dont understand this step,

OpenStudy (anonymous):

which part of it?

OpenStudy (unklerhaukus):

just this bit \[{\color\blue{\frac{d}{dt}\left(\frac{dy}{dx}\right)\frac{dt}{dx}}}\color\red=e^{-t} \frac{d}{dt}\left(e^{-t} \frac{dy}{dt}\right)\]

OpenStudy (anonymous):

oh u colored it... sorry... firstly we proved that \[ \frac{dy}{dx}=e^{-t} \frac{dy}{dt} \] and we know that \[ \frac{dt}{dx}=e^{-t} \]

OpenStudy (unklerhaukus):

hmm yeah

OpenStudy (unklerhaukus):

i think i get it now , thanks again \[x^2\frac{\text d^2y}{\text d{x^2}}+\alpha x\frac{\text dy}{\text dx}+by=f(x)\]\[\frac{\text dt}{\text dx}=e^{-t}\] \[\frac{\text dy}{\text dx}=\frac{\text dy}{\text dt} \frac{\text dt}{\text dx}=e^{-t} \frac{\text dy}{\text dt}\] \[\frac{\text d^2y}{\text dx^2}=\frac{\text d}{\text dx}\left(\frac{\text dy}{\text dx}\right)=\frac{\text d}{\text dx}\frac{\text dt}{\text dt}\left(e^{-t} \frac{\text dy}{\text dt}\right)=\frac{\text dt}{\text dx}\frac{\text d}{\text dt}\left(e^{-t} \frac{\text dy}{\text dt}\right)\]\[\qquad\qquad=e^{-t}\frac{\text d}{\text dt} \left(e^{-t} \frac{\text dy}{\text dt}\right)=e^{-2t}\frac{\text d^2 y}{\text dt^2}-e^{-2t}\frac{\text dy}{\text dt}\] \[e^{2t} \left(e^{-2t} \frac{\text d^2y}{\text dt^2}-e^{-2t}\frac{\text dy}{\text dt}\right)+\alpha e^{t} \left(e^{-t} \frac{\text dy}{\text dt}\right) +by=g(t) \] \[ \frac{\text d^2y}{\text dt^2}-\frac{\text dy}{\text dt}+\alpha \frac{\text dy}{\text dt}+by=g(t)\]\[ \frac{\text d^2y}{\text dt^2}+\left(\alpha-1\right)\frac{\text dy}{\text dt}+by=g(t)\]